Label the points O(0, 0), A(22, 4), and B(1, 1) on the graph
of f (x) 5 x 2. Determine the images of these points on the transformed function. Label the images Or, Ar, and Br.

f(x)=-(x-2)^2
1
Expert's answer
2020-12-06T10:27:58-0500

f(x) = −x4 +ax2 +5 a y = f(x) y. (0, 5). [ 2 marks]. 3b. Find . Markscheme ... Write down the -coordinates of these two points; ... The gradient of the tangent to the curve at a point P is.